An inductor-coil carries a steady-state current of 2.0 A when connected across an ideal battery an ideal battery of emf 4.0 V. If its inductance is 1.0 H, find the time constant of the circuit.

Grade:11

1 Answers

Navjyot Kalra
9 years ago
Sol. i = 2A, E = 4V, L = 1H R = E/i = 4/2 = 2 i = L/R = ½ = 0.5